3 If tan(x) = (in Quadrant 3), find 17 Give exact answers. role cos(22) sin n (1/2) = plz explain.
Answer #1

Alg tann = 3/4 H: 3+112 298 D 3 17 Cs5x = - 17 2 - IT 17-2-6- २११ sinx= -3 - 3 17.26 1238 sinx and sx are negative in hd quad